A 0.10M solution of NaOH ,is used to titrate 10.00ml of 0.10M CH3COOH.
Calculate the pH at equivalence point (3mks)
b) Sketch the titration curve in (a) above. (2mks)
c) Consider the titration of 25mL of 0.1M KCI with 0.050M AgNO3 using the Mohr method and given Ksp (AgCl) = 1.8 xl 0-10

Name the indicator used for the Mohr method. (1mk)
Use equations to explain how the end point is detected using this indicator.
1
Expert's answer
2020-04-08T13:31:03-0400
"CH_3COOH + NaOH = CH_3COONa + H_2O"

"pH = 7+0.5(pK_a+pC_b)"

"p = 7+0.5(-log_{10}(1.745*10^{-5}) -log_{10}(0.10)) = 7+2.88 = 9.88"

In the Mohr method, potassium chromate is used as an indicator.